OpenStudy (albert0898):

**Solve these systems of equations using addition to eliminate one of the variables. 14) 13a - b = 3 a + 3b = 11 21) a + 3b = 4 2a - b = 1

OpenStudy (amistre64):

we can multiply any equation by any value; and also subtract the equations (or add) them together. elimination is the process of multiplying the equations by a convient amount to get a variable that can be zeroed out

OpenStudy (amistre64):

once a variable is eliminated, the problem is resolved by solving for the remaing variable

OpenStudy (albert0898):

For question 14, I had gotten... 4a + 2b = 14 Now should I divide both side by 2 and solve from there?

OpenStudy (amistre64):

the problem there is you still have 2 variables so we havent made the problem easier. 13a - b = 3 a + 3b = 11 ^^ notice the bs here, if we multiply the top by 3 we can cancel these out 39a - 3b = 9 a + 3b = 11 ------------- 40a = 20 now its a cinch to solve for "a"

OpenStudy (amistre64):

we can just as easily multiply the bottom one by -13 to eliminate the "a" s and solve for b

OpenStudy (amistre64):

21 works in the same manner, but nicer since its got smaller values to play with

OpenStudy (albert0898):

But can I still divide both side by 2 and solve from there?

OpenStudy (amistre64):

lets try it for your idea and see if we can work it out 4a + 2b = 14 ; divide by 2 2a +b = 7 ; we still have no way of singling out a variable

OpenStudy (amistre64):

you might be able to substitute it into one of the others, but that kinda defeats the purpose of doing it this way to begin with :)

OpenStudy (albert0898):

But the direction state.. **Solve these systems of equations using 'addition' to eliminate one of the variables.

OpenStudy (amistre64):

right, and when you "added" you did not "eliminate" a variable

OpenStudy (amistre64):

first you multiply by a useful value to change the form of the eqautions into something that will eliminate when adding